OpenStudy (anonymous):

Which ratio is equivalent to 6 : 30?

HanAkoSolo (jamierox4ev3r):

just think about what a ratio expresses: 6/30

OpenStudy (anonymous):

this is multiple choice you need to attach the choices

OpenStudy (anonymous):

A. 1 : 6 B. 1 : 5 C. 1 : 3 D. 1 : 24

OpenStudy (anonymous):

you could multiply both numbers by 2 to get another ratio or you could decide them by the same number

OpenStudy (anonymous):

b

HanAkoSolo (jamierox4ev3r):

so if this ratio equals 6/30 ---> then you could simplify it to 3/15, etc.

OpenStudy (anonymous):

1:5 6/3=2 and 30/3=10 then 2/2=1 10/2-5 so 1 : 5

jimthompson5910 (jim_thompson5910):

To fully simplify the ratio, divide both parts by the GCF of 6 and 30.

OpenStudy (anonymous):

Try Multiplying Both By 2

HanAkoSolo (jamierox4ev3r):

what could 3/15 simplify to? @lee_actress … for future reference, @Cmonkeys32 , please do not provide answers without explaining the process a bit more thoroughly, thank you

OpenStudy (anonymous):

the answer is b because 6 goes into 30 5 times

OpenStudy (anonymous):

Thank you all!

OpenStudy (anonymous):

6*5=30

OpenStudy (anonymous):

Lol